### v1

# Unsupervised Learning for Quadratic Assignment

Abstract
We introduce PLUME search, a data-driven framework that enhances search efficiency in combinatorial optimization through unsupervised learning. Unlike supervised or reinforcement learning, PLUME search learns directly from problem instances using a permutation-based loss with a non-autoregressive approach. We evaluate its performance on the quadratic assignment problem, a fundamental NP-hard problem that encompasses various combinatorial optimization problems. Experimental results demonstrate that PLUME search consistently improves solution quality. Furthermore, we study the generalization behavior and show that the learned model generalizes across different densities and sizes.
1 Introduction
Combinatorial Optimization (CO) represents a central challenge in computer science and operations research, targeting optimal solutions within a large search space. CO encompasses numerous practical applications, including transportation logistics, production scheduling, network design, and resource allocation. The computational complexity of CO problems is typically NP-hard, making exact methods intractable for large instances. Researchers have thus developed approximation algorithms, heuristics, and hybrid approaches that balance solution quality with computational feasibility. These methods include simulated annealing, genetic algorithms, tabu search, and various problem-specific heuristics that can generate high-quality solutions within reasonable time budget Kirkpatrick et al. (1983); Holland (1992); Johnson and McGeoch (1997); Glover and Laguna (1998); Gomes and Selman (2001); Blum and Roli (2003).
1.1 Data-driven Combinatorial Optimization
Recently, data-driven methods have gained significant attention in addressing combinatorial optimization problems. Taking the Travelling Salesman Problem (TSP) as an example, researchers have explored both Supervised Learning (SL) and Reinforcement Learning (RL) methods. In SL, approaches such as pointer networks and graph neural networks attempt to learn mappings from problem instances to solutions by training on optimal or near-optimal tours Joshi et al. (2019); Vinyals et al. (2015). These models learn to imitate optimal or near-optimal solutions, leading to significant computational expense when building the training dataset. RL approaches frame TSP as a sequential decision-making problem where an agent learns to construct tours in a Markov decision process framework. While RL models have shown promise in small instances, these methods face significant challenges when scaling to larger problems. Furthermore, the sparse rewards and the high variance during training make it difficult for RL agents to learn effective policies Bello et al. (2016).
An alternative approach is Unsupervised Learning (UL), which avoids sequential decision making and does not require labelled data. In  Min et al. (2023), the authors propose a surrogate loss for the TSP objective by using a soft indicator matrix $\mathbb{T}$ to construct the Hamiltonian cycle, which is then optimized for minimum total distance. The $\mathbb{T}$ operator can be essentially interpreted as a soft permutation operator, as demonstrated in Min and Gomes (2023), which represents a rearrangement of nodes on the route $1\rightarrow 2\rightarrow 3\rightarrow\cdots\rightarrow n\rightarrow 1$, with $n$ representing the number of cities.
The TSP loss used in UL can be formally expressed in matrix notation. Essentially, we optimize:
$\mathcal{L}_{\text{TSP}}=\langle\mathbb{T}\mathbb{V}\mathbb{T}^{T},\mathbf{D}_{\text{TSP}}\rangle,$ (1)
where $\mathbb{V}$ represents a Hamiltonian cycle matrix encoding the route $1\rightarrow 2\rightarrow\cdots\rightarrow n\rightarrow 1$, $\mathbb{T}$ is an approximation of a hard permutation matrix $\mathbf{P},$ and $\mathbf{D}_{\text{TSP}}$ is the distance matrix with self-loop distances set to $\lambda$. $\mathbb{T}\mathbb{V}\mathbb{T}^{T}$ is a heat map that represents the probability that each edge belongs to the optimal solution, which guides the subsequent search process.
Since permutation operators are ubiquitous across many CO problems, and the application to TSP demonstrates their effectiveness, here, we extend this approach to a broader class of problems. Specifically, we propose Permutation-based Loss with Unsupervised Models for Efficient search (PLUME search), an unsupervised data-driven heuristic framework that leverages permutation-based learning to solve CO problems.
1.2 Quadratic Assignment Problem
The Quadratic Assignment Problem (QAP), essentially a permutation optimization problem, is an NP-hard problem with numerous applications across facility layout, scheduling, and computing systems. For example, in facility layout, QAP finds the optimal permutation of facilities minimizing total interaction costs based on inter-facility flows and inter-location distances. QAP’s applications also include manufacturing plant design, healthcare facility planning, VLSI circuit design, telecommunications network optimization, and resource scheduling Koopmans and Beckmann (1957); Lawler (1963).
Formally, QAP asks to assign $n$ facilities to $n$ locations while minimizing the total cost, which depends on facility interactions and location distances. A flow matrix $\mathbf{F}\in\mathbb{R}^{n\times n}$ captures the interaction cost between facility $i$ and facility $j$. Each location is represented by a matrix $\mathbf{X}\in\mathbb{R}^{n\times 2}$, where each row contains the coordinates of a location. The distance matrix $\mathbf{D}\in\mathbb{R}^{n\times n}$ is computed using the Euclidean distance $\mathbf{D}_{ij}=\|\mathbf{X}_{i}-\mathbf{X}_{j}\|_{2}$.
The objective is to find a permutation matrix $\mathbf{P}\in\{0,1\}^{n\times n}$ that maps facilities to locations while minimizing the total cost, given by $\min_{\sigma\in S_{n}}\sum_{i=1}^{n}\sum_{j=1}^{n}F_{ij}D_{\sigma(i),\sigma(j)}$, where $\sigma$ is a permutation of $\{1,\dots,n\}$ defining the assignment and $S_{n}$ denotes the set of all $n\times n$ permutation operators. This cost function can be equivalently written in matrix form as:
$\min_{\mathbf{P}\in S_{n}}\langle\mathbf{P}\mathbf{F}\mathbf{P}^{T},\mathbf{D}\rangle,$ (2)
where $\mathbf{P}$ is the permutation matrix representing the assignment.
1.3 Unsupervised Learning for QAP
In this paper, we use PLUME search to solve QAP. Following the TSP fashion in Min et al. (2023), we use neural networks (NNs) coupled with a Gumbel-Sinkhorn operator to construct a soft permutation matrix $\mathbb{T}$ that approximates a hard permutation matrix $\mathbf{P}$. We use the soft permutation $\mathbb{T}$ to guide the subsequent search process, as shown in Figure 1.
InputFeaturesNeuralNetworkPermutationFormulationOutputSearch
Figure 1: Overview of PLUME search framework: Input features are transformed through a neural network into a permutation formulation, then the output of the neural network guides the subsequent search process. This unified architecture allows PLUME to handle various combinatorial optimization problems by learning permutation operators.
PLUME search is a neural-guided heuristic that learns problem representations through NNs to directly guide the search. By integrating learned representations, PLUME search leverages UL to enhance search performance.
Here, for QAP, while TSP uses the heat map $\mathcal{H}=\mathbb{T}\mathbb{V}\mathbb{T}^{T}$ as a soft matrix (heat map) to guide the search, we directly decode a hard permutation matrix $\mathbf{P}$ from $\mathbb{T}$. This permutation matrix $\mathbf{P}$ represents an alignment in QAP and serves as an initialization to guide the subsequent search.
2 Tabu Search
We adopt Tabu search as the backbone of our PLUME search framework. Tabu Search (TS) is a method introduced by Glover Glover and Laguna (1998) that enhances local search methods by employing memory structures to navigate the solution space effectively. Unlike traditional hill-climbing algorithms, TS allows non-improving moves to escape local optima and uses adaptive memory to avoid cycling through previously visited solutions. The algorithm maintains a tabu list that prohibits certain moves, creating a dynamic balance between intensification and diversification strategies.
Tabu Search for Quadratic Assignment
TS has emerged as one of the most effective metaheuristics for addressing QAP instances Taillard (1991); Drezner (2003); James et al. (2009). The algorithm navigates the solution space through strategic move evaluations and maintaining memory structures to prevent cycling and encourage diversification. For QAP, TS typically begins with a random permutation as the initial solution and employs a swap-based neighborhood structure where adjacent solutions are generated by exchanging the assignments of two facilities. A key component of TS is the tabu list, which records recently visited solutions or solution attributes to prevent immediate revisiting. In QAP implementations, the tabu list typically tracks recently swapped facility pairs, prohibiting their re-exchange for a specified number of iterations—the tabu tenure. This memory structure forces the search to explore new regions of the solution space even when immediate improvements are not available, helping the algorithm escape local optima. The performance depends mainly on three parameters: neighbourhoodSize, which controls the sampling density from the complete swap neighbourhood at each iteration; evaluations, which establishes the maximum computational budget as measured by objective function calculations; and maxFails, which implements an adaptive early termination criterion that halts the search after a predefined number of consecutive non-improving iterations. Together, these parameters balance exploration breadth against computational efficiency, ensuring both effective solution space coverage and predictable runtime performance Glover (1989); Blum and Roli (2003); Battiti and Tecchiolli (1994); Misevicius (2005).
3 Model
We design NNs that preserve permutation equivalence while effectively encoding the underlying structure of QAP. Specifically, our NNs preserve permutation equivariance to ensure consistent handling of isomorphic problem instances. Furthermore, our model generalizes across variable problem sizes without modifying the NNs.
3.1 Permutation-Equivariant QAP Architecture
The NN’s structure is shown in Figure 2. Our permutation-equivariant architecture has two inputs: node positions $\mathbf{X}\in\mathbb{R}^{n\times 2}$ representing 2D coordinates and flow matrices $\mathbf{F}\in\mathbb{R}^{n\times n}$ representing pairwise interactions, where $n$ is the number of facilities. Our model consists of three main components: a position encoder, two flow feature encoders, and the main network. The position encoder maps facility coordinates through a shared Multi-Layer Perceptron (MLP) $\phi_{\text{pos}}$ to embed spatial information, the flow encoders process row-wise and column-wise interactions through pooling operations and MLPs:
$\mathbf{H}_{\text{pos}}=\phi_{\text{pos}}(\mathbf{X}),\quad\mathbf{H}_{\text{row}}=\psi_{\text{row}}\left(\sum_{j=1}^{N}\phi_{\text{row}}(\mathbf{F}_{ij})\right),\quad\mathbf{H}_{\text{col}}=\psi_{\text{col}}\left(\sum_{i=1}^{N}\phi_{\text{col}}(\mathbf{F}_{ij})\right),$ (3)
where $\phi_{\text{row}},\phi_{\text{col}}:\mathbb{R}\rightarrow\mathbb{R}^{d}$ and $\phi_{\text{pos}}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{d}$ are MLPs that transform input features into $d$-dimensional latent representations, and $\psi_{\text{row}},\psi_{\text{col}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ are MLPs that process the aggregated features to capture facility interactions. These features are concatenated and processed through $L$ permutation-equivariant MLP layers before the final projection:
$\mathbf{H}^{(0)}=[\mathbf{H}_{\text{pos}}\|\mathbf{H}_{\text{row}}\|\mathbf{H}_{\text{col}}],\quad\mathbf{H}^{(\ell+1)}=\text{MLP}^{(\ell)}(\mathbf{H}^{(\ell)}),\quad\mathbf{Y}=\Theta\mathbf{H}^{(L)},$ (4)
where $\mathbf{H}^{(0)}\in\mathbb{R}^{n\times 3d}$ contains the concatenated initial node features, $\mathbf{H}^{(\ell)}$ represents node embeddings at layer $\ell$, and $\text{MLP}^{(\ell)}$ denotes the $\ell$-th layer. These MLPs map from $3d$ to $d$ dimensions, while the final linear layer $\Theta\in\mathbb{R}^{d\times m}$ projects from $d$ to $m$ dimensions. $\mathbf{Y}\in\mathbb{R}^{n\times m}$ is the final output logits for assignment prediction. This architecture preserves permutation equivariance, that is, $f(\Pi\mathbf{X},\Pi\mathbf{F}\Pi^{T})=\Pi f(\mathbf{X},\mathbf{F})$ for any permutation matrix $\Pi\in S_{n}$, where $f$ represents the network function mapping inputs $\mathbf{X},\mathbf{F}$ to output $\mathbf{Y}$.
[…]
Figure 2: Neural network architecture for QAP. The model takes coordinates and flow matrix data and passes them into MLP encoders, the resulting features are then concatenated together. We then feed them into the main network and generate $\mathbf{Y}$ through an output layer.
3.2 Building Soft Permutation $\mathbb{T}$
Our model first generates logits which are transformed by a scaled hyperbolic tangent activation:
$\mathcal{F}=\alpha\tanh(\mathbf{Y}\mathbf{Y}^{T}),$ (5)
where $\alpha$ is a scaling parameter that controls the magnitude of the output. We then construct a soft permutation matrix $\mathbb{T}$ using the Gumbel-Sinkhorn operator:
$\mathbb{T}=\text{GS}(\frac{\mathcal{F}+\gamma\times\text{Gumbel noise}}{\tau},l),$ (6)
where GS denotes the Gumbel-Sinkhorn operator that builds a continuous relaxation of a permutation matrix. Here, $\gamma$ controls the scale of the Gumbel noise which adds stochasticity to the process, $\tau$ is the temperature parameter that controls the sharpness of the relaxation (lower values approximate discrete permutations more closely), and $l$ is the number of Sinkhorn normalization iterations.
During inference, we obtain a discrete permutation matrix $\mathbf{P}$ by applying the Hungarian algorithm to the scaled logits: $\mathbf{P}=\text{Hungarian}(-\frac{\mathcal{F}+\gamma\times\text{Gumbel noise}}{\tau})$. We use the CUDA implementation of the batched linear assignment solver for the Hungarian operator from  Karpukhin et al. (2024).
3.3 Invariance Property of Permutation Representation
The soft permutation matrix $\mathbb{T}$ is constructed through Equations 5 and 6 to preserve permutation equivariance. Let $\Pi\in S_{n}$ represent a random permutation on the nodes, the distance and flow matrices transform as $\mathbf{D}=\Pi\mathbf{D}_{0}\Pi^{T}$ and $\mathbf{F}=\Pi\mathbf{F}_{0}\Pi^{T}$, where $\mathbf{D}_{0}$ and $\mathbf{F}_{0}$ are the distance matrix and flow matrix before this random permutation respectively. Now, let $\mathbf{Y}_{0}$ denote the initial output, given that the network output transforms under permutation $\Pi$ as $\mathbf{Y}=\Pi\mathbf{Y}_{0}$, we then have $\mathbb{T}=\Pi\mathbb{T}_{0}\Pi^{T}$. Consequently, the objective function $\langle\mathbb{T}\mathbf{F}\mathbb{T}^{T},\mathbf{D}\rangle$ remains invariant. To be specific, $\forall~{}\Pi\in S_{n}$, we have:
$\langle\mathbb{T}\mathbf{F}\mathbb{T}^{T},\mathbf{D}\rangle$ $=\langle\Pi\mathbb{T}_{0}\Pi^{T}\Bigl{(}\Pi\mathbf{F}_{0}\Pi^{T}\Bigr{)}\Bigl{(}\Pi\mathbb{T}_{0}\Pi^{T}\Bigr{)}^{T},\Pi\mathbf{D}_{0}\Pi^{T}\rangle$ (7)
$=\langle\Pi\mathbb{T}_{0}\underbrace{\Pi^{T}\Pi}_{=I}\mathbf{F}_{0}\underbrace{\Pi^{T}\Pi}_{=I}\mathbb{T}^{T}_{0}\Pi^{T},\Pi\mathbf{D}_{0}\Pi^{T}\rangle=\langle\Pi\mathbb{T}_{0}\mathbf{F}_{0}\mathbb{T}^{T}_{0}\Pi^{T},\Pi\mathbf{D}_{0}\Pi^{T}\rangle$
$=\langle\Pi\Bigl{(}\mathbb{T}_{0}\mathbf{F}_{0}\mathbb{T}^{T}_{0}\Bigr{)}\Pi^{T},\Pi\mathbf{D}_{0}\Pi^{T}\rangle=\langle\mathbb{T}_{0}\mathbf{F}_{0}\mathbb{T}^{T}_{0},\mathbf{D}_{0}\rangle.$
This invariance property guarantees consistent solutions for isomorphic problem instances. Overall, Equations 5 and 6 preserve permutation symmetry while enabling gradient-based optimization, with the additional benefit of allowing the model to naturally generalize across different problem sizes. This generalization capability arises because the soft permutation matrix $\mathbb{T}\in\mathbb{R}^{n\times n}$ always matches the input graph size, independent of the model’s parameters. Consequently, our approach scales to problems of varying sizes without needing to modify the NN.
4 Results
4.1 Data Generation
We generate synthetic instances using an Erdős-Rényi (ER) graph model following Tan and Mu (2024). To build a QAP instance of size $n$, we generate a flow matrix $\mathbf{F}\in\mathbb{R}^{n\times n}$ and the location coordinates $\mathbf{X}=(\texttt{Uniform}(0,1),\texttt{Uniform}(0,1))$. We generate uniformly random weights in $[0,1]$ for the upper triangular portion of $\mathbf{F}$, then mirror these values to create a symmetric matrix. We apply an ER graph mask with edge probability $p$ to control the sparsity of connections between facilities.
Formally, for each QAP instance $i$:
$\mathbf{F}_{ij}$ =$=\begin{cases}\texttt{Uniform}(0,1)&\text{if }\text{rand}()<p\text{ and }i\neq j\\ 0&\text{otherwise}\end{cases},\quad\forall i<j$ (8)
$\mathbf{F}_{ji}$ $=\mathbf{F}_{ij},\quad\forall i<j.$ (9)
We build datasets with varying problem sizes $n\in\{100,200\}$ and graph densities $p\in\{0.1,0.2,\ldots,0.9\}$. For each configuration, we generate 30,000 instances for training, 5,000 for validation and 5,000 for test, respectively. We run experiments using an NVIDIA H100 GPU and an Intel Xeon Gold 6154 CPU.
4.2 Training
We optimize our NNs to minimize the QAP objective:
$\langle\mathbb{T}\mathbf{F}\mathbb{T}^{T},\mathbf{D}\rangle,$ (10)
with the model’s hidden dimension set to either $32$ or $64$. The number of layers $l$ is set to either $2$ or $4$. For the Gumbel-Softmax operator used in Equation 5, we set $\tau\in\{1,2,3,4,5\}$ and $l\in\{10,20,30,40,50\}$. The noise scale $\gamma$ is set to either $0.001$ or $0.01$. The tanh scale is set to $\alpha=40$. We use the Adam optimizer with a learning rate of $3\times 10^{-5}$ and train for $100$ epochs. We then select the model with the best validation performance to evaluate on the test set. For problem size $100$, we set $m=100$, and for problem size $200$, we set $m=200$.
4.3 Method
We train our model on each problem size $n$ and each graph density $p$. After training the model, we then validate it and select the best parameters before testing. We implement a PLUME search in a straightforward way. As mentioned, the model outputs the soft permutation matrix $\mathbb{T}$, and we decode the hard permutation matrix $\mathbf{P}$ from $\mathbb{T}$. $\mathbf{P}$ corresponds to a learned assignment. We then start the tabu search for QAP using this learned assignment as the initial solution.
Table 1: Comparison between the average costs of UL predicted solutions and random solutions. The gap value indicates the average percentage improvement of the predicted solution over the random solution’s cost. Inference time denotes the average total time (running NN inference+Hungarian) required to build the UL predicted solutions.
$n=100$ $n=200$
$p$ UL random Gap (%) Inference (ms) UL random Gap (%) Inference (ms)
$0.1$ 230.086 257.877 10.7768 $2.80944\times 10^{-2}$ 958.645 1037.01 7.55682 $1.91477\times 10^{-1}$
$0.2$ 478.116 515.381 7.23057 $3.14968\times 10^{-2}$ 1968.86 2074.41 5.08819 $1.82030\times 10^{-1}$
$0.3$ 730.449 773.459 5.56073 $2.98188\times 10^{-2}$ 2987.00 3110.27 3.96332 $1.96343\times 10^{-1}$
$0.4$ 984.907 1031.92 4.55588 $3.15542\times 10^{-2}$ 4009.58 4144.96 3.26614 $1.93695\times 10^{-1}$
$0.5$ 1239.22 1288.90 3.85445 $2.83198\times 10^{-2}$ 5040.71 5184.32 2.77008 $1.92888\times 10^{-1}$
$0.6$ 1496.41 1547.62 3.30895 $2.83606\times 10^{-2}$ 6072.39 6219.82 2.37033 $1.86270\times 10^{-1}$
$0.7$ 1754.76 1805.83 2.82806 $3.20354\times 10^{-2}$ 7112.34 7257.90 2.00554 $1.91129\times 10^{-1}$
$0.8$ 2025.43 2064.49 1.89199 $2.93868\times 10^{-2}$ 8154.81 8298.41 1.73045 $1.93583\times 10^{-1}$
$0.9$ 2278.93 2320.06 1.77280 $2.78900\times 10^{-2}$ 9199.26 9336.99 1.47510 $1.89588\times 10^{-1}$
Mean 1246.48 1289.50 3.33618 $2.96619\times 10^{-2}$ 5055.96 5184.90 2.48684 $1.90778\times 10^{-1}$
4.4 Effectiveness of UL-Based Initialization
Before diving into the PLUME search’s final results, we first check whether the learned assignment improves the solution quality without any subsequent search. We directly compare the quality of solutions using learned assignments versus random assignments in the initialization stage.
Table 1 demonstrates the cost of using UL-predicted solutions compared to random initialization. We define the gap as:
$1-\frac{\langle\mathbf{P}\mathbf{F}\mathbf{P}^{T},\mathbf{D}\rangle}{\langle\mathbf{P}_{random}\mathbf{F}\mathbf{P}^{T}_{random},\mathbf{D}\rangle},$
where $\mathbf{P}$ is the learned assignment and $\mathbf{P}_{random}$ is a random assignment. The gap percentage shows consistent improvement across all problem densities, with more significant gains observed in sparser problems. For $n=100$ with density $p=0.1$, the UL approach achieves a 10.7770% improvement over random initialization, while for $n=200$, it yields a 7.55672% improvement. As problem density increases, the gap narrows but remains positive, indicating that our approach maintains its advantage even for denser problems.
[…]
TS($\mu$,$\kappa$,$\omega$) $\text{TS}(1k,25,25)$ $\text{TS}(1k,100,25)$ $\text{TS}(1k,25,100)$ $\text{TS}(5k,100,100)$ $\text{TS}(10k,100,100)$
$p$ UL random UL random UL random UL random UL random
$0.1$ 190.129 198.931 205.433 224.427 190.154 198.888 175.675 181.297 165.349 167.250
$0.2$ 422.156 434.907 443.269 469.997 422.089 434.915 478.116 410.615 388.068 390.903
$0.3$ 663.959 679.353 688.723 720.382 663.984 679.369 640.422 650.792 623.827 627.494
$0.4$ 911.248 928.427 938.500 973.599 911.228 928.329 885.413 896.848 867.245 871.310
$0.5$ 1160.92 1179.51 1189.85 1227.30 1161.02 1179.49 1133.85 1146.22 1114.71 1119.23
$0.6$ 1415.86 1435.01 1445.24 1484.27 1415.84 1435.06 1387.92 1400.88 1368.54 1373.03
$0.7$ 1673.38 1692.78 1703.09 1742.15 1673.31 1692.70 1645.29 1658.21 1625.78 1630.24
$0.8$ 1939.07 1953.37 1972.66 2001.89 1939.01 1953.46 1910.17 1919.61 1889.13 1892.27
$0.9$ 2198.54 2213.92 2229.21 2260.26 2198.55 2213.89 2171.44 2181.57 2152.00 2155.33
Average 1175.03 1190.69 1201.78 1233.81 1175.021 1190.68 1150.26 1160.67 1132.74 1136.34
[…]
TS($\mu$,$\kappa$,$\omega$) $\text{TS}(1k,25,25)$ $\text{TS}(1k,100,25)$ $\text{TS}(1k,25,100)$ $\text{TS}(5k,100,100)$ $\text{TS}(10k,100,100)$
$p$ UL random UL random UL random UL random UL random
$0.1$ 874.188 918.224 915.225 981.426 874.146 918.342 836.399 871.458 798.478 816.504
$0.2$ 1851.52 1913.14 1908.32 1998.87 1851.68 1912.98 1799.74 1849.18 1748.06 1773.71
$0.3$ 2848.72 2921.98 2987.00 3022.16 2848.77 2921.71 2788.05 2847.12 2727.83 2758.39
$0.4$ 3856.81 3937.86 3929.77 4048.26 3856.62 3937.92 3790.02 3856.20 3724.59 3758.61
$0.5$ 4878.52 4965.70 4955.69 5082.12 4878.88 4965.68 4807.95 4879.02 4738.94 4776.09
$0.6$ 5905.06 5994.81 5984.58 6114.63 5905.33 5994.91 5832.49 5905.84 5761.67 5800.11
$0.7$ 6943.11 7032.42 7023.56 7152.33 6943.24 7032.42 6869.62 6942.84 6798.78 6836.88
$0.8$ 7988.29 8076.76 8067.27 8194.76 7988.28 8077.04 7916.56 7988.74 7846.75 7884.92
$0.9$ 9039.82 9125.34 9115.28 9237.99 9040.02 9124.94 8971.13 9041.09 8904.87 8941.46
Average 4909.56 4987.36 4979.41 5092.50 4909.66 4987.33 4845.77 4909.06 4783.33 4816.30
4.5 PLUME Search for QAP
[…]
Notably, the relative improvement from using PLUME search is more significant for larger problem sizes, with $n=200$ showing consistently higher percentage improvements than $n=100$ at comparable densities. This trend holds for tabu search with different parameters. For example, with parameters $\mu=1,000$, $\kappa=25$, and $\omega=25$, the tabu search with random initialization achieves $1190.69$ on average for $n=100$, while PLUME search achieves $1175.03$, yielding a 1.31520% improvement. At $n=200$, PLUME search achieves $4909.56$ on average while random initialized tabu search achieves $4987.36$, resulting in a more substantial 1.55994% improvement. The most intensive tabu search configuration, TS$(10k,100,100)$, provides the most comprehensive exploration of the solution space and thus yields the best solution quality across all initialization methods. At this parameter setting, PLUME search further improves solution quality. For graphs with $n=100$, PLUME search achieves an average solution cost of $1132.74$ compared to $1136.34$ for random initialization, yielding a 0.316807% improvement. While this may appear modest, it is important to note that as search parameters increase, the final solutions converge closer to optimality, leaving less room for improvement. The fact that PLUME search maintains its advantage even in this setting underscores the quality of its initialization. For larger graphs with $n=200$ vertices, the benefit becomes more pronounced. PLUME search achieves an average cost of $4783.33$ compared to $4816.30$ for random initialization, yielding a 0.684550% improvement. This increasing advantage with problem size suggests that PLUME search scales well to larger problems.
Table 4: Average Time comparison of tabu search with different search parameters on QAP instances with $n=100$, in ms (×$\times 10^{-3}$ s).
TS($\mu$,$\kappa$,$\omega$) $\text{TS}(1k,25,25)$ $\text{TS}(1k,100,25)$ $\text{TS}(1k,25,100)$ $\text{TS}(5k,100,100)$ $\text{TS}(10k,100,100)$
$p$ UL random UL random UL random UL random UL random
$0.1$ 1.55014 1.47786 2.65438 2.01335 2.02096 2.32776 3.96913 3.86332 6.89681 6.40729
$0.2$ 1.40809 1.72910 2.16947 2.11550 2.19710 2.13621 4.38935 4.10793 6.19193 6.92497
$0.3$ 2.23220 2.30824 2.32095 2.20613 2.03746 2.12630 4.06415 4.28922 6.34255 6.50180
$0.4$ 2.72354 2.91837 2.04465 2.24711 1.91479 2.45331 4.05188 4.01972 6.73373 6.90415
$0.5$ 2.57762 1.60781 2.13888 2.35198 1.97659 2.44347 4.20058 4.24511 6.87188 6.36137
$0.6$ 1.90425 1.71858 2.26572 2.21770 2.34254 1.99306 3.86804 3.87752 6.63121 6.72185
$0.7$ 1.60661 1.57086 2.11504 2.34658 2.11745 2.13709 3.66044 3.91046 6.60751 6.77780
$0.8$ 1.91665 2.06804 2.02229 2.31294 2.28833 1.97817 3.77077 3.66541 6.81585 6.88984
$0.9$ 2.42608 2.11719 2.11124 2.00790 2.31411 2.31984 3.67512 4.25899 6.28867 6.47949
Mean 2.03835 1.94623 2.20474 2.20213 2.13437 2.21280 3.96105 4.02641 6.59779 6.66317
Table 5: Average Time comparison of tabu search with different search parameters on QAP instances with $n=200$, in ms (×$\times 10^{-3}$ s).
TS($\mu$,$\kappa$,$\omega$) $\text{TS}(1k,25,25)$ $\text{TS}(1k,100,25)$ $\text{TS}(1k,25,100)$ $\text{TS}(5k,100,100)$ $\text{TS}(10k,100,100)$
$p$ UL random UL random UL random UL random UL random
$0.1$ 2.42600 2.29224 2.26313 2.38713 2.49597 2.70196 6.85827 6.82996 11.8025 11.8930
$0.2$ 2.29030 2.33664 2.38857 2.28357 2.78697 2.75465 6.62369 6.64074 11.9798 11.8780
$0.3$ 2.34508 2.32658 2.27949 2.36495 2.65163 2.77468 6.49103 6.87804 11.9024 11.7376
$0.4$ 2.18291 2.57973 2.33208 2.16072 2.85288 2.80369 6.79153 6.76758 12.0805 11.9279
$0.5$ 2.34563 2.22272 2.54632 2.50635 2.74198 2.82831 6.59106 6.78405 12.0350 11.9095
$0.6$ 2.54630 2.29580 2.33564 2.56085 2.54337 2.97725 6.65846 6.81566 12.0519 11.6976
$0.7$ 2.27941 2.27283 2.54844 2.33566 2.48008 3.13651 7.27975 7.16379 12.1370 11.8553
$0.8$ 2.35310 2.39341 2.60996 2.48153 2.37177 2.01584 7.05032 6.80656 11.6703 11.8593
$0.9$ 2.18062 2.51180 2.41696 2.68309 2.06515 2.72407 7.16032 7.26328 11.6522 11.5438
Mean 2.32771 2.35908 2.41340 2.41821 2.55442 2.74633 6.83383 6.88330 11.92351 11.81135
Tables 4 and 5 show the execution time comparison (TS runtime) between random and UL-based initialization approaches for tabu search. The inference time for generating UL-predicted solutions is negligible: $<3\times 10^{-2}$ ms for $n=100$ and $<2\times 10^{-1}$ ms for $n=200$ (shown in Table 1). When accounting for both inference and execution time, the total computational cost of UL-based initialization remains comparable to random initialization across all parameter configurations. Overall, our results show that PLUME search effectively improves the efficiency of the tabu search algorithm, enabling the discovery of better solutions within a similar computational budget.
4.6 Compared with Other Data-driven Methods
We compare with recent work by  Tan and Mu (2024), where the authors use RL and test only on QAP instances up to size 100. We run their model on $n=100$ and $p=0.7$, using their model, we observe an average cost of $1644.37$ with an average time of $\approx 150$ ms, our PLUME TS($10010k,100,$ 100) achieves better performance with a cost of $1625.78$. Notably, our methods are substantially faster, with average time costs of just $7$ ms. It should be noted that although Tan and Mu (2024) also uses tabu search as a benchmark; they employ a Python implementation, which is not as computationally efficient as our C++ implementation of tabu search. We didn’t fine-tune the RL method and fine-tuning may yield better performance. However, the gaps are so dramatically large that even with optimization, the RL approach would remain substantially inferior.
5 Generalization
Table 6: Generalization across densities for the model trained on $n=100$, $p=0.7$ and test on $n=100$ with different densities. Random initialization shows costs using random initial assignments. UL-based initialization shows costs using assignments predicted by our neural network. TS$(\mu,\kappa,\omega)$ shows costs after running tabu search from random initialization. PLUME $\text{TS}(\mu,\kappa,\omega)$ shows costs after running tabu search from UL-predicted initialization.
Density $0.1$ $0.2$ $0.3$ $0.4$ $0.5$ $0.6$ $0.8$ $0.9$
Random Initialization 257.877 515.381 773.459 1031.92 1288.90 1547.62 2064.49 2320.06
UL-based Initialization 258.090 515.928 774.288 1032.26 1282.81 1528.17 2049.10 2320.75
$\text{TS}(1k,25,25)$ 198.931 434.907 679.353 928.427 1179.51 1435.01 1953.37 2213.92
PLUME $\text{TS}(1k,25,25)$ 198.991 434.811 679.407 928.081 1177.45 1428.55 1947.81 2213.91
$\text{TS}(10k,100,100)$ 167.250 390.903 627.494 871.310 1119.23 1373.03 1892.27 2155.33
PLUME $\text{TS}(10k,100,100)$ 167.296 390.828 627.470 871.144 1118.73 1371.62 1891.02 2155.40
Table 7: Generalization across densities for the model trained on $n=200$, $p=0.7$ and test on $n=200$ with different densities.
Density $0.1$ $0.2$ $0.3$ $0.4$ $0.5$ $0.6$ $0.8$ $0.9$
Random Initialization 1037.01 2074.41 3110.27 4144.96 5184.32 6219.82 8298.41 9336.99
UL-based Initialization 1036.78 1968.86 3110.34 4199.06 5040.71 6072.39 8284.74 9358.93
$\text{TS}(1k,25,25)$ 918.299 1913.04 2921.80 3937.95 4965.65 5994.98 8076.76 9125.18
PLUME $\text{TS}(1k,25,25)$ 918.402 1851.55 2921.83 3965.58 4878.55 5905.10 8069.67 9137.63
$\text{TS}(10k,100,100)$ 816.661 1773.70 2758.83 3758.61 4776.40 5800.17 7884.66 8941.58
PLUME $\text{TS}(10k,100,100)$ 816.572 1748.18 2758.66 3768.02 4738.89 5761.62 7882.97 8946.10
5.1 Cross-density Generalization
We further study how the model generalizes across different densities, as shown in Table 6 and 7. Using models trained on a specific edge density ($p=0.7$), our results show the best performance on nearby densities ($p=0.6$ and $p=0.8$). Model performance gradually declines as we move further from the training density, with more significant drops at the extreme values ($p=0.1$ and $p=0.9$). This suggests the model captures transferable structural patterns that work best within a reasonable proximity to its training conditions. This effect is more pronounced with TS$(1k,25,25)$ compared to configurations with more extensive search parameters. On $n=100$, with TS$(1k,25,25)$, the performance difference between PLUME and random initialization at $p=0.6$ is 0.450171%, while with TS$(10k,100,100)$, the difference narrows to 0.102693%.
Table 8: Cross-size generalization: Model trained on $n=200$ instances and tested on $n=100$ instances (both with density $p=0.7$). Random Initialization shows cost from random assignments; UL-based Initialization shows cost from UL model. TS$(\mu,\kappa,\omega)$ shows costs after running tabu search from random initialization. PLUME $\text{TS}(\mu,\kappa,\omega)$ shows costs after running tabu search from UL-predicted initialization.
Random Initialization UL-based Initialization TS $(1k,25,25)$ PLUME TS $(1k,25,25)$ TS $(10k,100,100)$ PLUME TS $(10k,100,100)$
1805.83 1795.57 1692.78 1688.43 1630.24 1629.54
Table 9: Cross-size generalization: Model trained on $n=100$ instances and tested on $n=200$ instances (both with density $p=0.7$).
Random Initialization UL-based Initialization TS $(1k,25,25)$ PLUME TS $(1k,25,25)$ TS $(10k,100,100)$ PLUME TS $(10k,100,100)$
7257.90 7233.49 7023.566 7015.50 6836.88 6830.74
5.2 Cross-size Generalization
Our model naturally generalizes across problem sizes due to its permutation-equivariant design, where the soft permutation matrix $\mathbb{T}\in\mathbb{R}^{n\times n}$ automatically adapts to match input dimensions. Experiments show that a model trained on $n=100$ effectively generalizes to $n=200$ problems and vice versa, while consistently outperforming random initialization, as shown in Table 8 and Table 9. For instance, when a model trained on $n=200$ is applied to $n=100$ problems, UL initialization achieves a solution cost of 1795.57 compared to 1805.83 for random initialization, and when used with TS$(1k,25,25)$, PLUME TS reaches 1688.43 versus 1692.78 for standard TS with random initialization.
6 Conclusion
In this paper, we propose PLUME search, a framework to enhance combinatorial optimization through UL. By leveraging permutation-based loss, we demonstrate that neural networks can effectively learn QAP directly from instances instead of using SL/RL. Our experimental results suggest that UL can generate high-quality initial solutions that significantly outperform random initialization and these better starting points consistently lead to better final solutions after tabu search. This is achieved by a simple NN structure with negligible computational overhead. Our approach demonstrates strong generalization across varying problem densities and sizes.
PLUME search takes a different approach from traditional heuristics and other SL/RL methods, offering a complementary method that integrates seamlessly with existing frameworks rather than competing with them. Future work might explore extending this approach to other permutation-based combinatorial problems, developing more sophisticated neural architectures to capture problem structure, and investigating the theoretical foundations of UL for combinatorial optimization.
7 Acknowledgement
This project is partially supported by the Eric and Wendy Schmidt AI in Science Postdoctoral Fellowship, a Schmidt Futures program; the National Science Foundation (NSF) and the National Institute of Food and Agriculture (NIFA); the Air Force Office of Scientific Research (AFOSR); the Department of Energy; and the Toyota Research Institute (TRI).